OpenStudy (anonymous):

Find the standard formula for the hyperbola with vertices at (-7, 0) and (7, 0) and foci at (-9, 0 ) and (9, 0).

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

can someone help me!

OpenStudy (shadowfiend):

So it looks like this hyperbola is centered around (0, 0), which will make things easier.

OpenStudy (shadowfiend):

It also looks like it opens left and right. Thus, the standard form equation should be: \[\frac{x^2}{a^2} - \frac{y^2}{b^2}\] That means we just have to find a and b.

OpenStudy (shadowfiend):

To find a and b, we need to find the distance from the center to the foci. In this case, that's easy: it's 9. That means that we have: \[9^2 = a^2 + b^2\] \[81 = a^2 + b^2\]

OpenStudy (anonymous):

thanks so much shadowfiend!! i really appreciate your help

OpenStudy (shadowfiend):

We also have the vertices of the hyperbola. We know the vertices for a hyperbola centered at (0, 0) are: (a, 0) and (-a, 0). Since the vertices here are (7, 0) and (-7, 0), we know a = 7. Thus, we have: \[\begin{align} 81 &= 7^2 + b^2\\ 81 &= 49 + b^2\\ 81 - 49 &= b^2\\ 32 &= b^2 \end{align}\] We don't need b itself, just \(b^2\), for the equation of the hyperbola. We plugin \(a^2\) and \(b^2\) and get: \[\frac{x^2}{49} - \frac{y^2}{32} = 1\] Does that all make sense?

OpenStudy (anonymous):

yeah that makes sense thanks alot... how do I reward you with a medal?

OpenStudy (shadowfiend):

Click the `good answer' button on the right :)

OpenStudy (anonymous):

I'm not seeing a good answer button... sorry I'm new to this site

OpenStudy (shadowfiend):

Damn, really?

OpenStudy (anonymous):

could you point it out

OpenStudy (shadowfiend):

Yeah, one sec.

OpenStudy (shadowfiend):

Attached a screenshot showing it. Let me know if you still can't find it :)

OpenStudy (amistre64):

looks like shadow is hulking out :)

OpenStudy (anonymous):

thats so strange it doesn't show for me...

OpenStudy (shadowfiend):

That's definitely a problem. tree2193, could you post a screenshot of your screen? I definitely need to see if we can get that fixed.

OpenStudy (anonymous):

I WANT YOU TO GET YOUR REWARD MEDAL!!!

OpenStudy (anonymous):

yeah sure how do I do that?

OpenStudy (shadowfiend):

Are you running Windows?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

hit f5 to refresh the browser and you should see a 'good answer' beside his name

OpenStudy (shadowfiend):

Ok, cool. There should be a button somewhere with the text `prt scrn' or `print screen' or `print scrn' or some variant like that. Hit Alt+that button.

OpenStudy (shadowfiend):

Then open Paint and hit Ctrl-V to paste it, and save that file and post it here. amistre -- have you run into the good answer button not showing up before?

OpenStudy (amistre64):

yep, mostly in internet explorer

OpenStudy (shadowfiend):

Ah. Well that's a doozy :p We will definitely have to take a look at why that's happening.

OpenStudy (anonymous):

ok I got it!

OpenStudy (amistre64):

f5 tends to cure it tho

OpenStudy (anonymous):

I just had to press f5

OpenStudy (shadowfiend):

Ok, cool. Thanks to both of you. We'll see if we can figure out why IE's being silly this time. And thanks for the medal tree2193!

OpenStudy (anonymous):

I have a similar problem if you have time

OpenStudy (shadowfiend):

I probably don't... But post it, and if no one else gets to it in a little while, I'll swing by and check and see what I can do.

OpenStudy (anonymous):

ok thanks!

OpenStudy (shadowfiend):

No problem!